Find the general solution of 4 -8 2 cosht 12 -6 (cosht(+32 sinh) cosht + 2 sinh NI ( -8)*+
Answer #1

Q find the find the general general som of 2 cosht 어 + de x ( )* 6 cosht ta sinht sol : 4 -8 let A- -b also of Athen eigenvave (1 = to da=2 (i) eigen vector corresponding (A-25) v=0 2 -8 :):)-13) » - 8 Q 2x, 8 X 2 an 2 E R uan 2 = » if is then eigenA cash e + B sinh t 3 Xplt) = c cash t + D simht diff w. r.do A simht + B cosht 4 x (4) c simht + D cash t then from o 2 coshA simht + B cosht => c sinht to cusht un - 8c) cosht + (4B – 80) simht (2A-6c) conht + (2B-60) sinnt 2 cosht + cosht + a sinhthen general som of of 0 e is X(t) Xc (1) + Xolt) = -ut at + C₂ e 0 xlt) eet = C, e 2 sinht + simht which is required soon.